# Mixture of Contexts for Long Video Generation

### v1

@@ Abstract
Long video generation is fundamentally a long-context memory problem: models must retain and retrieve salient events across long range without collapsing or drifting. However, scaling diffusion transformers (DiTs) to generate long-context videos is fundamentally limited by the quadratic cost of self-attention, which makes memory and computation intractable and difficult to optimize for long sequences. We recast long-context video generation as an internal information retrieval task and propose a simple, learnable sparse attention routing module, Mixture of Contexts (MoC), as an effective long-term memory retrieval engine. In MoC, each query dynamically selects a few informative chunks plus mandatory anchors (caption, local windows) to attend to, with causal routing that prevents loop closures. As we scale the data and gradually sparsify the routing, the model allocates compute to salient history, preserving identities, actions, and scenes over minutes of content. Efficiency follows as a byproduct of retrieval (near-linear scaling), which enables practical training and synthesis, and the emergence of memory and consistency at the scale of minutes.
@@ 1 Introduction
Video generation has emerged as a central problem in generative modeling, powering content creation, simulation for autonomous systems, and interactive storytelling. Recent Transformer-based diffusion models can synthesize increasingly realistic clips by modeling complex space–time dependencies; yet, pushing them to minute- or hour-long horizons exposes a deeper challenge: long-term memory. Models must retain and retrieve salient events across extended timelines without drift, collapse, or loss of identity. Dense self-attention becomes computationally prohibitive as sequences grow, and moreover, the core difficulty is not merely computational, but learning to selectively recall the right context at the right time.
@@
In this work, we reformulate long-context video generation as an internal information retrieval process, where each token dynamically accesses only the most relevant context through learnable sparse attention routing. To realize this, we propose an adaptive Mixture of Contexts (MoC) framework that learns to route each query to the most relevant segments of the video sequence, instead of relying on uniform or static sparse attention or a fixed selection strategy. Specifically, MoC partitions the multi-modal token stream into content-aligned chunks along frames, shots, and captions, then lets each query select only a few relevant chunks via a parameter-free yet trainable top-$k$ router. Two mandatory anchors: cross-modal links to all text tokens and intra-shot local window links are activated to stabilize local fidelity while reserving routing capacity for genuinely long-range recall. A causal routing mask is additionally applied to prevent pathological loop closures by enforcing a directed acyclic interaction graph, improving roll-out robustness over minute-scale sequences. For efficient implementation, the selected key tokens are directly processed by the flash-attention kernel, which supports variable sequence lengths and high throughput. During training, we progressively adjust the granularity of chunks and the selectivity of the routing mechanism, resulting in a gradual sparsification that encourages the model to focus on the most informative context as training progresses.
@@ 2 Related Work
Long Video Generation. Existing video generation models [23, 36, 13, 47, 15, 1, 17, 5, 6, 12, 51] are mostly limited to a few seconds. Several recent frameworks specifically target longer video generation using autoregressive models that operate on frames, chunks, or segments, such as CausVid [48]. While these frameworks extend generation capabilities, they often grapple with error accumulation [38] inherent in sequential prediction or face uncertain computational scaling to longer durations. To mitigate these issues, RollingDiffusion [31] and Diffusion Forcing [3] inject controlled noise into the historical context and train the model to denoise it, increasing robustness to compounding errors. MAGI-1 [32] and SkyReels-V2 [4] scale up these ideas by employing autoregressive denoising, aiming for potentially longer durations. An orthogonal strategy is to distill the entire past into a constant-size latent. TTTVideo [7] uses a learnable MLP to encode the context during inference, while FramePack [55] encodes arbitrarily many frames into a fixed vector for next-frame prediction. FramePack [55] also proposes early planning of future frames to mitigate the error accumulation issue. This is similar to using keyframes or anchor frames [16, 39, 26, 59, 19, 45, 46, 43], where certain frames are predefined and the video generation model only does an interpolation sampling job. These methods extend video generation to the one-minute range but still face a hard ceiling on maintaining long-context coherence going forward, as they rely on lossy compression of the contexts. The work most closely related to ours is Long-Context Tuning [14] (LCT), which starts from a single-shot DiT and expands its context window to a scene comprising up to eight shots ($\approx$8s, $\sim 2.3{\times}10^{4}$ tokens each). LCT [14] keeps the attention mechanism dense: all text and video tokens inside the enlarged window attend to one another after being positioned with an interleaved 3D RoPE. While this design elegantly re-uses the pretrained weights and yields impressive multi-shot coherence, it inherits the quadratic cost of full self-attention – FLOPs and memory scale with $(8L_{\text{shot}})^{2}$.
@@
Sparse Attention for Video Generation. Sparse attention leverages the observation that attention matrices are often sparse (many scores are near zero) and computes attention only for a subset of important token pairs, a natural fit for video generation given spatiotemporal redundancy. Training-free pruners include SparseVideoGen [41], which profiles heads that dynamically specialize into spatial vs. temporal and selects a per-head pattern, and STA [57], which exploits localized 3D windows by operating tile-by-tile over FlashAttention-friendly blocks [9, 8]. Universal filters such as SpargeAttn/SageAttention [54, 53, 52] combine selective token compression with a softmax-aware pass to skip parts of $QK^{T}$/PVPV, and AdaSpa [42] proposes a “blockified” dynamic pattern with Fused LSE-Cached Search that reuses sparse indices across denoising steps. Jenga [58] uses training-free block-wise attention carving plus progressive resolution. Beyond these post-hoc pruners, recent trainable or structured designs include VMoBA [40], which learns a mixture-of-block scheme with layer-wise partitions and global/thresholded block selection for VDMs. VSA [56] proposes a hardware-efficient coarse-to-fine sparse kernel that replaces full attention at both training and inference. Radial Attention [25] instead, uses a static $\mathcal{O}(n\log n)$ mask derived from spatiotemporal energy-decay that enables longer generations with near-dense quality. While these advances substantially reduce costs and accelerate video generation, most methods either prune emergent dense maps or impose fixed sparsity priors, focusing on accelerating the generation of short videos. By contrast, our Mixture of Contexts learns deliberate, end-to-end routing of context sources and focuses on long context memory/consistency.
@@
Context Learning in Visual Generation. A complementary line of work treats context—past frames, states, or reference images—as a first-class signal for learning and control. For video world models, where action and camera position signals are available, WORLDMEM [44] augments simulators with an external memory bank of frames and states and retrieves relevant entries via Field-of-View (FoV) overlapping to preserve long-term scene consistency. A similar work, Context-as-Memory [49], targets interactive long videos, explicitly retrieving a small set of historical frames as conditions for each step to sustain scene consistency, also via FoV overlapping to select the relevant frames. Back to the image space, IC-LoRA [20] demonstrates that DiTs already exhibit in-context abilities and proposes concatenating reference images with lightweight task-specific LoRA [18] to adapt across tasks with few samples. DSD [2] turns in-context generation into paired supervision via self-distillation: curate image grids with a VLM, then fine-tune a text+image-to-image model. OminiControl [35] offers a parameter-efficient, unified framework for image-conditioned control in DiTs, enabling broad conditioned tasks without auxiliary modules. Recent open-sourced models, such as FLUX-Context [24], concatenate text and images to unify in-context image generation and editing, with improved consistency. These works demonstrate that, given a sufficiently large training scale, routing and in-context learning are very powerful in extracting useful information from contexts. Our Mixture of Contexts follows this routine, and proposes to learn to route among multiple context sources end-to-end, enabling deliberate selection and composition of contextual signals rather than relying solely on fixed retrieval or a single conditioning pathway.
@@ 3.2 Attention Chunking and Routing
Fixed Cross-Modal Selection. In addition to dynamically routed visual chunks, we explicitly require every visual query token to attend to all text tokens in the sequence since text tokens, while typically constituting less than 1% of all tokens, encode the most semantically informative signals—specifying global style, character identities, and key actions. The computational overhead is negligible, yet the benefits are substantial: anchoring generation to the prompt significantly reduces prompt-drift errors and prevents the fading of rare attribute words during long video roll-outs. Furthermore, this hard cross-modal link facilitates joint gradient propagation into both text and visual embeddings, tightening their shared latent space and markedly improving editability in downstream tasks such as text-guided video editing.
Fixed Intra-Shot Selection. Long videos naturally exhibit a strict hierarchical structure, with frames nested within shots and shots within scenes. To leverage this, we explicitly enforce the intra-shot connections in the attention mechanism, ensuring that each token always attends to its belonging shots—capturing object trajectories, lighting continuity, and other predictive cues. This design allows the Mixture of Contexts (MoC) framework to allocate its sparse attention budget to genuinely long-range dependencies, rather than redundantly modeling local context. Enforcing such connections offers several benefits: it guards against semantic discontinuities at scene cuts where adjacent tokens may become unrelated; it guarantees that every attention matrix contains at least one well-conditioned block; and it provides a contiguous, memory-efficient fallback path even under aggressive adaptive pruning. This strategy is particularly effective when fine-tuning pretrained video generation models, as it preserves the fidelity of each shot from the outset and enables the model to gradually learn to align broader contextual information during training.
@@ 3.3 Computation Efficiency
Base Model 0.9380 0.9623 0.9816 0.6875 0.5200 0.6345 0% $1.9\times 10^{10}$
Ours 0.9398 0.9670 0.9851 0.7500 0.5547 0.6396 83% $\mathbf{4.1\times 10^{9}}$
Table 1: Single-shot video generation quantitative comparison. We report VBench [21] metrics and computation efficiency metrics. Our method is on par with or better than the base model for all VBench [21] metrics despite aggressive sparsification (83%).
[…]
Consistency Consistency Smoothness Degree Quality Quality
@@ 4 Experiments
We mainly conduct our experiment on two tasks: shot-level text-to-video generation and scene-level text-to-video generation with multiple shot cuts.
@@
Base model. We build our model on a long-context video generator, LCT [14], which supports both single-shot and multi-shot video generation. LCT adapts a 3B-parameter MMDiT [10] architecture that was trained on a mixture of image, single-shot, and multi-shot videos at their native resolutions and durations. The model’s full self-attention is expanded from per-shot scope to a scene-level context window of up to eight shots (roughly 8 seconds, 22k tokens each), using an interleaved 3D RoPE [34] to give every shot distinct absolute coordinates while preserving the relative layout of text and video tokens. We initialize our model weights from pretrained LCT [14] and replace its attention module with our MoC, then fine-tune using the identical training scheme as LCT [14].
[…]
@@
Evaluation Metrics. For both short single-shot video generation and long multi-shot video generation, we follow prior work [48, 55] and evaluate on the popular VBench [21, 22] benchmark. Specifically, Subject Consistency and Background Consistency indicate how faithfully the primary subject and background from the input image are preserved throughout the video, Motion Smoothness evaluates the fluidity of movement (lack of jitter or abrupt transitions), and Dynamic Degree measures the extent of motion in the video (encouraging the generation of dynamic content rather than static scenes). We also report Aesthetic Quality and Image Quality for single-shot videos, to quantify each frame’s visual appeal and technical quality. In addition, we report computational metrics such as sparsity, FLOPs, and inference speedup compared with Flash Attention [9, 8].
@@
Quantitative Results. Tab. 1 and Tab. 2 present a quantitative comparison between our content-aligned Mixture of Contexts (MoC) model and dense attention baselines on both short single-shot clips and long multi-shot scenes. For short single-shot videos (6k tokens), despite the aggressive sparsification, our method matches or surpasses the dense baseline across all VBench metrics. This demonstrates that directing computational resources toward the most relevant chunks not only reduces FLOPs but also enables the model to maintain character fidelity and scene coherence better. However, for such short sequences, the additional overhead from index gathering and pooling outweighs the computational savings, resulting in a slower end-to-end pipeline. In contrast, for long multi-shot videos (180k tokens), MoC exhibits clear computational advantages. By discarding 85% of the context, our approach achieves a $2.2\times$ speedup. Furthermore, it substantially enhances the performance of our model, particularly in terms of motion diversity, as evidenced by an increase in Dynamic-Degree from 0.46 to 0.56, while maintaining Motion-Smoothness. Although this increased motion budget leads to a slight reduction in appearance fidelity, all quality metrics remain high. Collectively, these results validate the core premise of our approach: learned, structure-aware sparsity reallocates computation from redundant frames to salient visual events, delivering significant efficiency gains without compromising (and in some cases improving) perceptual quality.
Qualitative Results. We present qualitative comparisons in Figs. 3 and 4 for single-shot and multi-shot video generation tasks, respectively.
Figure 3: Single-shot video generation qualitative comparison. Our results are on par, if not better than, our base model despite aggressive sparsification.
@@ 5 Conclusion
Adaptive Mixture of Contexts (MoC) demonstrates that learnable sparse attention routing can function as a powerful, data-driven memory retrieval engine. Our work is arguably the first to show that by scaling up training data with an efficient and learnable sparse routing mechanism, a model can develop a sophisticated method for long-term recall. This approach achieves minute-scale memory at a cost comparable to short-video generation. Critically, this capability emerges without explicit heuristics like 3D priors or fixed-rule selection; the model learns entirely from data which historical context is salient. Because the routing is learned and the implementation is fast during inference, MoC provides a blueprint for the next generation of scalable, controllable, and responsible long-video generative models. It proves that removing the quadratic attention bottleneck is not just an efficiency gain but a direct path to unlocking emergent, long-term memory in video generation.

### v2

@@ Abstract
Long-context video generation is fundamentally a memory problem: models must retain and retrieve salient events across long range without collapsing or drifting. However, scaling diffusion transformers (DiTs) to generate long-context videos is fundamentally limited by the quadratic cost of self-attention, which makes memory and computation intractable and difficult to optimize for long sequences. We recast long-context video generation as an internal information retrieval task and propose a simple, learnable sparse attention routing module, Mixture of Contexts (MoC), as an effective long-term memory retrieval engine. In MoC, each query dynamically selects a few informative chunks plus mandatory anchors (caption, local windows) to attend to, with causal routing that prevents loop closures. As we scale the data and gradually sparsify the routing, the model allocates compute to salient history, preserving identities, actions, and scenes over minutes of content. Efficiency follows as a byproduct of retrieval (near-linear scaling), which enables practical training and synthesis, and the emergence of memory and consistency at the scale of minutes.
@@ 1 Introduction
Video generation has emerged as a central problem in generative modeling, powering content creation, simulation for autonomous systems, and interactive storytelling. Recent Transformer‑based diffusion models can synthesize increasingly realistic clips by modeling complex space–time dependencies; yet, pushing them to minute- or hour-long horizons exposes a deeper challenge: long-term memory. Models must retain and retrieve salient events across extended timelines without drift, collapse, or loss of identity. Dense self-attention becomes computationally prohibitive as sequences grow, and moreover, the core difficulty is not merely computational, but learning to selectively recall the right context at the right time.
@@
In this work, we reformulate long-context video generation as an internal information retrieval process, where each token dynamically accesses only the most relevant context through learnable sparse attention routing. To realize this, we propose an adaptive Mixture of Contexts (MoC) framework that learns to route each query to the most relevant segments of the video sequence, instead of relying on uniform or static sparse attention or a fixed selection strategy. Specifically, MoC partitions the multi‑modal token stream into content‑aligned chunks along frames, shots, and captions, then lets each query select only a few relevant chunks via a parameter-free yet trainable top‑$k$ router. Two mandatory anchors: cross‑modal links to all text tokens and intra‑shot local window links are activated to stabilize local fidelity while reserving routing capacity for genuinely long‑range recall. A causal routing mask is additionally applied to prevent pathological loop closures by enforcing a directed acyclic interaction graph, improving roll-out robustness over minute‑scale sequences. For efficient implementation, the selected key tokens are directly processed by the flash-attention kernel, which supports variable sequence lengths and high throughput. During training, we progressively adjust the granularity of chunks and the selectivity of the routing mechanism, resulting in a gradual sparsification that encourages the model to focus on the most informative context as training progresses.
@@ 2 Related Work
Long Video Generation. Existing video generation models [23, 37, 13, 49, 15, 1, 17, 5, 6, 12, 53] are mostly limited to a few seconds. Several recent frameworks specifically target longer video generation using autoregressive models that operate on frames, chunks, or segments, such as CausVid [50]. While these frameworks extend generation capabilities, they often grapple with error accumulation [39] inherent in sequential prediction or face uncertain computational scaling to longer durations. To mitigate these issues, RollingDiffusion [32] and Diffusion Forcing [3] inject controlled noise into the historical context and train the model to denoise it, increasing robustness to compounding errors. MAGI-1 [33] and SkyReels-V2 [4] scale up these ideas by employing autoregressive denoising, aiming for potentially longer durations. An orthogonal strategy is to distill the entire past into a constant-size latent. TTTVideo [7] and LaCT[60] use a learnable MLP to encode the context during inference, while FramePack [57] encodes arbitrarily many frames into a fixed vector for next-frame prediction. FramePack [57] also proposes early planning of future frames to mitigate the error accumulation issue. This is similar to using keyframes or anchor frames [16, 40, 27, 62, 19, 47, 48, 45], where certain frames are predefined and the video generation model only does an interpolation sampling job. These methods extend video generation to the one-minute range but still face a hard ceiling on maintaining long-context coherence going forward, as they rely on lossy compression of the contexts. The work most closely related to ours is Long-Context Tuning [14] (LCT), which starts from a single-shot DiT and expands its context window to a scene comprising up to eight shots ($\approx$8s, $\sim 2.3{\times}10^{4}$ tokens each). LCT [14] keeps the attention mechanism dense: all text and video tokens inside the enlarged window attend to one another after being positioned with an interleaved 3D RoPE. While this design elegantly re-uses the pretrained weights and yields impressive multi-shot coherence, it inherits the quadratic cost of full self-attention – FLOPs and memory scale with $(8L_{\text{shot}})^{2}$.
@@
Sparse Attention for Video Generation. Sparse attention leverages the observation that attention matrices are often sparse (many scores are near zero) and computes attention only for a subset of important token pairs, a natural fit for video generation given spatiotemporal redundancy. Training‑free pruners include SparseVideoGen [42], which profiles heads that dynamically specialize into spatial vs. temporal and selects a per‑head pattern, and STA [59], which exploits localized 3D windows by operating tile‑by‑tile over FlashAttention‑friendly blocks [9, 8]. Universal filters such as SpargeAttn/SageAttention [56, 55, 54] combine selective token compression with a softmax‑aware pass to skip parts of $QK^{T}$/PVPV, and AdaSpa [43] proposes a “blockified” dynamic pattern with Fused LSE‑Cached Search that reuses sparse indices across denoising steps. Jenga [61] uses training‑free block‑wise attention carving plus progressive resolution. Beyond these post-hoc pruners, recent trainable or structured designs include VMoBA [41], which learns a mixture-of-block scheme with layer-wise partitions and global/thresholded block selection for VDMs. VSA [58] proposes a hardware‑efficient coarse‑to‑fine sparse kernel that replaces full attention at both training and inference. Radial Attention [26] instead, uses a static $\mathcal{O}(n\log n)$ mask derived from spatiotemporal energy‑decay that enables longer generations with near‑dense quality. While these advances substantially reduce costs and accelerate video generation, most methods either prune emergent dense maps or impose fixed sparsity priors, focusing on accelerating the generation of short videos. By contrast, our Mixture of Contexts learns deliberate, end‑to‑end routing of context sources and focuses on long context memory/consistency, with acceleration as a byproduct of sparsity.
@@
Context Learning in Visual Generation. A complementary line of work treats context—past frames, states, or reference images—as a first‑class signal for learning and control. For video world models, where action and camera position signals are available, WORLDMEM [46] augments simulators with an external memory bank of frames and states and retrieves relevant entries via Field-of-View (FoV) overlapping to preserve long‑term scene consistency. A similar work, Context‑as‑Memory [51], targets interactive long videos, explicitly retrieving a small set of historical frames as conditions for each step to sustain scene consistency, also via FoV overlapping to select the relevant frames. Concurrently, VMem [25] uses a surfel‑indexed, occlusion‑aware memory to retrieve relevant views and maintain consistency under re-visits. Back to the image space, IC‑LoRA [20] demonstrates that DiTs already exhibit in-context abilities and proposes concatenating reference images with lightweight task-specific LoRA [18] to adapt across tasks with few samples. DSD [2] turns in‑context generation into paired supervision via self‑distillation: curate image grids with a VLM, then fine‑tune a text+image‑to‑image model. OminiControl [36] offers a parameter‑efficient, unified framework for image‑conditioned control in DiTs, enabling broad conditioned tasks without auxiliary modules. Recent open-sourced models, such as FLUX‑Context [24], concatenate text and images to unify in-context image generation and editing, with improved consistency. These works demonstrate that, given a sufficiently large training scale, routing and in-context learning are very powerful in extracting useful information from contexts. Our Mixture of Contexts follows this routine, and proposes to learn to route among multiple context sources end‑to‑end, enabling deliberate selection and composition of contextual signals rather than relying solely on fixed retrieval or a single conditioning pathway.
@@ 3.2 Attention Chunking and Routing
Fixed Cross-Modal Selection as Attention Sink. In addition to dynamically routed visual chunks, we explicitly require every visual query token to attend to all text tokens in the sequence. This design mirrors a naive use of “sink tokens [44]”: a small, persistent set of tokens that every query can attend to, which (i) provides a low-entropy, semantically meaningful anchor for the attention distribution, (ii) guarantees at least one well-conditioned dense block in each attention matrix, and (iii) creates a global gradient highway. We directly use the text tokens, as they typically constitute less than 1% of all tokens, while encoding the most semantically informative signals—specifying global style, character identities, and key actions. The computational overhead is negligible, yet the benefits are substantial: anchoring generation to the prompt significantly reduces prompt-drift errors and prevents the fading of rare attribute words during long video roll-outs. Furthermore, this hard cross-modal link facilitates joint gradient propagation into both text and visual embeddings, tightening their shared latent space and markedly improving editability in downstream tasks such as text-guided video editing.
Fixed Intra-Shot Selection as Local Window. Long videos naturally exhibit a strict hierarchical structure, with frames nested within shots and shots within scenes. To leverage this, we explicitly enforce the intra-shot connections in the attention mechanism, ensuring that each token always attends to its belonging shots—capturing object trajectories, lighting continuity, and other predictive cues. This design allows the Mixture of Contexts (MoC) framework to allocate its sparse attention budget to genuinely long-range dependencies, rather than redundantly modeling local context. Enforcing such connections offers several benefits: it guards against semantic discontinuities at scene cuts where adjacent tokens may become unrelated; it guarantees that every attention matrix contains at least one well-conditioned block; and it provides a contiguous, memory-efficient fallback path even under aggressive adaptive pruning. This strategy is particularly effective when fine-tuning pretrained video generation models, as it preserves the fidelity of each shot from the outset and enables the model to gradually learn to align broader contextual information during training.
@@ 4 Experiments
We conduct our main experiment on long scene-level text-to-video generation with multiple shot cuts, a significant use case in AIGC video generation.
@@
Base model. We build our model on a long-context video generator, LCT [14], which is the only available architecture that supports long, multi-shot video generation for general scenes. LCT adapts a 3B-parameter MMDiT [10] architecture that was trained on a mixture of images, single-shot, and multi-shot videos at their native resolutions and durations. The model’s full self-attention is expanded from per-shot scope to a scene-level context window of up to eight shots (roughly 8 seconds, 22k tokens each), using an interleaved 3D RoPE [35] to give every shot distinct absolute coordinates while preserving the relative layout of text and video tokens. We initialize our model weights from pretrained LCT [14] and replace its attention module with our MoC, then fine-tune using the identical training scheme as LCT [14].
Baselines. We compare MoC with the base model LCT [14], which uses dense attention. For these experiments, we test on 8-shot sequences, where each shot is an 8-second 480p video with 12 FPS. This yields roughly 180k tokens per 64-second scene.
@@
Evaluation Metrics. We follow prior work [50, 57] and evaluate on the popular VBench [21, 22] benchmark. Specifically, Subject Consistency and Background Consistency indicate how faithfully the primary subject and background from the input image are preserved throughout the video, Motion Smoothness evaluates the fluidity of movement (lack of jitter or abrupt transitions), and Dynamic Degree measures the extent of motion in the video (encouraging the generation of dynamic content rather than static scenes). We also report Aesthetic Quality and Image Quality to quantify each frame’s visual appeal and technical quality. In addition, we report computational metrics such as sparsity, FLOPs, and inference speedup compared with Flash Attention [9, 8].
@@
Quantitative Results. Tab. 1 presents a quantitative comparison between our content-aligned Mixture of Contexts (MoC) model and dense attention baseline on minute-long multi-shot scenes. MoC exhibits clear computational advantages. By discarding 85% of the context, our approach achieves a $2.2\times$ speedup. Furthermore, it substantially enhances the performance of our model, particularly in terms of motion diversity, as evidenced by an increase in Dynamic-Degree from 0.46 to 0.56, while maintaining Motion-Smoothness. Although this increased motion budget leads to a slight reduction in appearance fidelity, all quality metrics remain high. Collectively, these results validate the core premise of our approach: learned, structure-aware sparsity reallocates computation from redundant frames to salient visual events, delivering significant efficiency gains without compromising (and in many cases improving) perceptual quality.
Qualitative Results. We present qualitative comparisons in Fig. 3.
@@ 5 Conclusion
Adaptive Mixture of Contexts (MoC) demonstrates that learnable sparse attention routing can function as a powerful, data-driven memory retrieval engine. Our work is arguably the first to show that by scaling up training data with an efficient and learnable sparse routing mechanism, a model can develop a sophisticated method for long-term recall. This approach achieves minute-scale memory at a cost comparable to short-video generation. Critically, this capability emerges without explicit heuristics like 3D priors or Field-of-View (FoV) selection; the model learns entirely from data which historical context is salient. Because the routing is learned and the implementation is fast during inference, MoC provides a blueprint for the next generation of scalable, controllable, and responsible long-video generative models. It proves that removing the quadratic attention bottleneck is not just an efficiency gain but a direct path to unlocking emergent, long-term memory in video generation.